# To the Noise and Back: Diffusion for Shared Autonomy

### v3

@@ II-B Diffusion model guidance for distribution transformation
3:$\tilde{\bm{a}}^{h}_{t,k_{\textrm{sw}}}\leftarrow\sqrt{\bar{\alpha}_{k_{\textrm{sw}}}}\bm{a}^{h}_{t}+\sqrt{1-\bar{\alpha}_{k_{\textrm{sw}}}}\bm{\epsilon}$triangleright Forward process
4:$\bm{a}^{s}_{t,k_{\textrm{sw}}}\leftarrow\tilde{\bm{a}}^{h}_{t,k_{\textrm{sw}}}$
@@
7:    $\bm{a}^{s}_{t,k-1}\leftarrow\bm{\mu}_{\theta^{*}}(\bm{a}^{s}_{t,k},k\mid\bm{s}_{t})+\sigma_{k}\mathbf{z}$triangleright Reverse process
@@ II-C Distribution transformation for shared autonomy
Figure 3: We evaluate our algorithm in the context of four shared autonomy environments including a LABEL:sub@fig:maze_env_render 2D Control task in which an agent navigates to one of two different goals, LABEL:sub@fig:ll-lander_env_render Lunar Lander that tasks a drone with landing at a designated location, LABEL:sub@fig:ll-reacher_env_render a Lunar Reacher variant in which the objective is to reach a designated region in the environment, and LABEL:sub@fig:bp_env_render Block Pushing, in which the objective is to use a robot arm to push an object into one of two different goal regions.
@@ III Experiments
Pilot
w/o Copilot
w/ Copilot
w/o Copilot
w/ Copilot
w/o Copilot
w/ Copilot
w/o Copilot
w/ Copilot
Noisy
$20.67\pm 4.50$
$\bm{68.00\pm 5.35}$
$\hphantom{0}28.33\pm 2.62$
$\hphantom{0}\bm{7.67\pm 2.87}$
$14.33\pm 2.49$
$\bm{45.33\pm 3.30}$
$\hphantom{0}77.33\pm 3.09$
$\bm{38.00\pm 2.94}$
Laggy
$21.33\pm 2.05$
$\bm{75.00\pm 3.56}$
$\hphantom{0}76.67\pm 2.49$
$\hphantom{0}\bm{9.67\pm 3.86}$
$30.67\pm 5.56$
$\bm{55.33\pm 6.13}$
$\hphantom{0}69.33\pm 5.56$
$\bm{31.33\pm 2.87}$
Zero
$0\hphantom{0}0.00\pm 0$ 0.00
$27.00\pm 0.82$
$100.00\pm 0.00$
$19.00\pm 2.94$
$0\hphantom{0}0.00\pm 0$ 0.00
$19.67\pm 2.62$
$100.00\pm 0.00$
$58.33\pm 3.09$
Random
$0\hphantom{0}0.00\pm 0$ 0.00
$25.00\pm 4.32$
$100.00\pm 0.00$
$19.33\pm 5.25$
$\hphantom{0}4.33\pm 1.89$
$22.33\pm 2.87$
$\hphantom{0}95.33\pm 1.70$
$59.00\pm 0.82$
Expert
$77.67\pm 2.62$
$78.67\pm 2.87$
$\hphantom{0}12.33\pm 0.94$
$\hphantom{0}8.00\pm 1.63$
$49.33\pm 4.78$
$55.00\pm 2.16$
$\hphantom{0}44.00\pm 3.56$
$31.67\pm 2.87$
@@ III-A Continuous control domains
[$\gamma=0.0$]Refer to caption [$\gamma=0.1$]Refer to caption [$\gamma=0.2$]Refer to caption
[$\gamma=0.4$]Refer to caption [$\gamma=0.6$]Refer to caption [$\gamma=0.8$]Refer to caption
@@ III-A2 Lunar Lander
Figure 5: Success (higher is better), floating, and crash (lower is better) rates for Lunar Lander (top) and Lunar reacher (bottom) with LABEL:sub@fig:ll-land-assisted-noisy LABEL:sub@fig:ll-reach-assisted-noisy noisy and LABEL:sub@fig:ll-land-assisted-laggy LABEL:sub@fig:ll-reach-assisted-laggy laggy pilots. The dashed blue line denotes the success rate of an expert policy, while the dotted blue line denotes the success rate of our model with full-diffusion ($\gamma=1.0$).
@@ III-A4 Block Pushing
Figure 6: The rates of block pushing task to terminate at correct goal, wrong goal or timeout with LABEL:sub@fig:ll-reach-assisted-noisy noisy and LABEL:sub@fig:ll-reach-assisted-laggy laggy pilot. The dashed blue line denotes the rate at which an expert policy reaches the correct goal, while the dotted blue line denotes the same rate for our model with full-diffusion ($\gamma=1.0$).
@@ III-B Training
To understand the pilot’s intent, our copilot relies on a state-conditioned diffusion model $p_{\theta}(a_{t}|s_{t})$. This diffusion model is trained as follows. First, we collect expert demonstrations, each containing a sequence of goal-embedded state-action pairs for each domain, where the goal is randomly sampled at each episode. We then use the resulting demonstrations to train a state-conditioned diffusion model using the DDPM loss (Eqn. II-A). The details of the model architecture are described in Appendix A. As we hide a goal from our copilot, we remove goal locations from each observation prior to training.
@@
[…]
@@ III-D Various pilots with our copilot
Pilot
w/o Copilot
w/ Copilot
w/o Copilot
w/ Copilot
Noisy
$62.33\pm 2.49$
$\bm{75.67\pm 2.05}$
$\hphantom{0}37.33\pm 2.05$
$\bm{20.00\pm 3.27}$
Laggy
$42.00\pm 4.90$
$\bm{74.33\pm 2.87}$
$\hphantom{0}58.00\pm 4.90$
$\bm{21.33\pm 4.50}$
Zero
$0\hphantom{0}0.00\pm 0$ 0.00
$40.67\pm 2.62$
$100.00\pm 0.00$
$\hphantom{0}6.67\pm 1.70$
Random
$0\hphantom{0}0.00\pm 0$ 0.00
$16.33\pm 3.09$
$100.00\pm 0.00$
$68.00\pm 1.63$
Expert
$99.00\pm 0.82$
$94.67\pm 2.05$
$\hphantom{00}1.00\pm 0.82$
$\hphantom{0}5.33\pm 2.05$
@@ III-F Real Human User Experiments
We conducted a set of human user experiments involving the Lunar Lander and Lunar Reacher continuous control domains. For the experiments, we recruited $17$ participants ($9$ who identified as male, and $8$ as female, with an average age of $25$). [^3] None of the participants were co-authors or otherwise involved in this research. We asked each user to interact with one of two different co-pilots within each episode, one corresponding to direct teleoperation (i.e., no assistance) and the other being our diffusion-based shared autonomy assistant. The identity of the co-pilot was not disclosed to the participant. At the beginning of each experiment, we allowed the user to practice for $10$ episodes with each co-pilot. In the subsequent testing phase, the user controlled the system for another $10$ episodes with each of the co-pilots. We conducted this experiment for both Lunar Lander and Lunar Reacher and emphasize that the location of the goal (the landing pad for Lunar Lander and goal region for Lunar Reacher) was chosen randomly from the continuous space of goals for each episode.
@@
Success Rate $\uparrow$
Crash Rate $\downarrow$
Float Rate
Success Rate $\uparrow$
Crash Rate $\downarrow$
Float Rate
w/o Copilot
[…]
$98.24$
$0.00$
$17.06$
$82.94$
$0.00$
w/ Copilot
$32.35$
$61.18$
$6.47$
$34.71$
$65.29$
$0.00$
@@ III-G Real-Robot Experiments
We recruited $14$ participants ($10$ who identified as male, $4$ who identified as female, average age of $27.7$) [^4] None of the participants were co-authors or otherwise involved in this research. for human user experiments with the real robot.
@@
success rate $\uparrow$
episode length $\downarrow$
w/o Copilot
$0.89$
$209.23\pm 83.71$
w/ Copilot
$1.00$
$143.66\pm 68.42$
@@ Appendix A Model architecture and hyperparameters
We generate the set of demonstrations used to train our diffusion model by collecting $100$k transitions for each goal in the 2D Control environment using the scripted expert policy provided in the original implementation [18], $3$M transitions for each goal in the Block Pushing environment using the pretrained expert, and $10$M transitions for Lunar Lander and Lunar Reacher using the corresponding pretrained expert. When we collect expert demonstrations, we filter out episodes that do not succeed, effectively making the quality of the demonstrations better than the original expert. This explains why the assisted expert often performs better than the expert (e.g., Table I—Lunar Reacher).
@@ B-A Lunar Lander and Lunar Reacher
[$p_{\text{noise}}=0.2$]Refer to caption [$p_{\text{noise}}=0.3$]Refer to caption [$p_{\text{noise}}=0.4$]Refer to caption [$p_{\text{noise}}=0.5$]Refer to caption
[$p_{\text{noise}}=0.6$]Refer to caption [$p_{\text{noise}}=0.7$]Refer to caption [$p_{\text{noise}}=0.8$]Refer to caption [$p_{\text{noise}}=0.9$]Refer to caption
@@
[$p_{\text{laggy}}=0.2$]Refer to caption [$p_{\text{laggy}}=0.3$]Refer to caption [$p_{\text{laggy}}=0.4$]Refer to caption [$p_{\text{laggy}}=0.5$]Refer to caption  [$p_{\text{laggy}}=0.6$]Refer to caption [$p_{\text{laggy}}=0.7$]Refer to caption [$p_{\text{laggy}}=0.8$]Refer to caption [$p_{\text{laggy}}=0.9$]Refer to caption
@@
[$p_{\text{noisy}}=0.2$]Refer to caption [$p_{\text{noisy}}=0.3$]Refer to caption [$p_{\text{noisy}}=0.4$]Refer to caption [$p_{\text{noisy}}=0.5$]Refer to caption  [$p_{\text{noisy}}=0.6$]Refer to caption [$p_{\text{noisy}}=0.7$]Refer to caption [$p_{\text{noisy}}=0.8$]Refer to caption [$p_{\text{noisy}}=0.9$]Refer to caption
@@
[$p_{\text{laggy}}=0.2$]Refer to caption [$p_{\text{laggy}}=0.3$]Refer to caption [$p_{\text{laggy}}=0.4$]Refer to caption [$p_{\text{laggy}}=0.5$]Refer to caption  [$p_{\text{laggy}}=0.6$]Refer to caption [$p_{\text{laggy}}=0.7$]Refer to caption [$p_{\text{laggy}}=0.8$]Refer to caption [$p_{\text{laggy}}=0.9$]Refer to caption
@@
Pilot
w/o Copilot ($\gamma=0.0$)
w/ Copilot (ours) ($\gamma=0.4$)
w/ Copilot ($\gamma=1.0$)
w/o Copilot ($\gamma=0.0$)
w/ Copilot (ours) ($\gamma=0.4$)
w/ Copilot ($\gamma=1.0$)
@@
Noisy
$20.67\pm 4.50$
$\bm{68.00\pm 5.35}$
$29.00\pm 2.94$
$\hphantom{0}28.33\pm 2.62$
$\bm{7.67\pm 2.87}$
$18.33\pm 4.92$
Laggy
$21.33\pm 2.05$
$\bm{75.00\pm 3.56}$
$26.33\pm 4.78$
$\hphantom{0}76.67\pm 2.49$
$\bm{9.67\pm 3.86}$
$15.67\pm 0.94$
Zero
$0\hphantom{0}0.00\pm 0$ 0.00
$27.00\pm 0.82$
$28.00\pm 2.16$
$100.00\pm 0.00$
$19.00\pm 2.94$
$13.33\pm 2.49$
Random
$0\hphantom{0}0.00\pm 0$ 0.00
$25.00\pm 4.32$
$27.67\pm 1.70$
$100.00\pm 0.00$
$19.33\pm 5.25$
$15.67\pm 1.25$
Expert
$77.67\pm 2.62$
$78.67\pm 2.87$
$29.33\pm 6.18$
$\hphantom{0}12.33\pm 0.94$
$\hphantom{0}8.00\pm 1.63$
$15.33\pm 1.70$
@@
Noisy
$14.33\pm 2.49$
$\bm{45.33\pm 3.30}$
$30.67\pm 4.11$
$\hphantom{0}77.33\pm 3.09$
$\bm{38.00\pm 2.94}$
$47.00\pm 3.56$
Laggy
$30.67\pm 5.56$
$\bm{55.33\pm 6.13}$
$30.67\pm 2.62$
$\hphantom{0}69.33\pm 5.56$
$\bm{31.33\pm 2.87}$
$45.67\pm 0.94$
Zero
$0\hphantom{0}0.00\pm 0$ 0.00
$19.67\pm 2.62$
$31.00\pm 6.98$
$100.00\pm 0.00$
$58.33\pm 3.09$
$45.33\pm 5.56$
Random
$\hphantom{0}4.33\pm 1.89$
$22.33\pm 2.87$
$30.00\pm 4.08$
$\hphantom{0}95.33\pm 1.70$
$59.00\pm 0.82$
$45.00\pm 2.16$
Expert
$49.33\pm 4.78$
$55.00\pm 2.16$
$28.67\pm 4.50$
$\hphantom{0}44.00\pm 3.56$
$31.67\pm 2.87$
$50.00\pm 1.41$

### v4

@@ II-B Diffusion model guidance for distribution transformation
[…]
@@
7:  $\bm{a}^{s}_{t,k-1}\leftarrow\bm{\mu}_{\theta^{*}}(\bm{a}^{s}_{t,k},k\mid\bm{s}_{t})+\sigma_{k}\mathbf{z}$ $\triangleright$ Reverse process
@@ II-C Distribution transformation for shared autonomy
Figure 3: We evaluate our algorithm in the context of four shared autonomy environments including a LABEL:sub@fig:maze_env_render 2D Control task in which an agent navigates to one of two different goals, LABEL:sub@fig:$l$-lander_env_render Lunar Lander that tasks a drone with landing at a designated location, LABEL:sub@fig:$l$-reacher_env_render a Lunar Reacher variant in which the objective is to reach a designated region in the environment, and LABEL:sub@fig:bp_env_render Block Pushing, in which the objective is to use a robot arm to push an object into one of two different goal regions.
@@ III Experiments
Pilot w/o Copilot w/ Copilot w/o Copilot w/ Copilot w/o Copilot w/ Copilot w/o Copilot w/ Copilot
Noisy $20.67\pm 4.50$ $\bm{68.00\pm 5.35}$ $\hphantom{0}28.33\pm 2.62$ $\hphantom{0}\bm{7.67\pm 2.87}$ $14.33\pm 2.49$ $\bm{45.33\pm 3.30}$ $\hphantom{0}77.33\pm 3.09$ $\bm{38.00\pm 2.94}$
Laggy $21.33\pm 2.05$ $\bm{75.00\pm 3.56}$ $\hphantom{0}76.67\pm 2.49$ $\hphantom{0}\bm{9.67\pm 3.86}$ $30.67\pm 5.56$ $\bm{55.33\pm 6.13}$ $\hphantom{0}69.33\pm 5.56$ $\bm{31.33\pm 2.87}$
Zero $\hphantom{0}0.00\pm 0.00$ $27.00\pm 0.82$ $100.00\pm 0.00$ $19.00\pm 2.94$ $\hphantom{0}0.00\pm 0.00$ $19.67\pm 2.62$ $100.00\pm 0.00$ $58.33\pm 3.09$
Random $\hphantom{0}0.00\pm 0.00$ $25.00\pm 4.32$ $100.00\pm 0.00$ $19.33\pm 5.25$ $\hphantom{0}4.33\pm 1.89$ $22.33\pm 2.87$ $\hphantom{0}95.33\pm 1.70$ $59.00\pm 0.82$
Expert $77.67\pm 2.62$ $78.67\pm 2.87$ $\hphantom{0}12.33\pm 0.94$ $\hphantom{0}8.00\pm 1.63$ $49.33\pm 4.78$ $55.00\pm 2.16$ $\hphantom{0}44.00\pm 3.56$ $31.67\pm 2.87$
@@ III-A Continuous control domains
$\gamma=0.0$]Refer to caption $\gamma=0.1$]Refer to caption $\gamma=0.2$]Refer to caption
$\gamma=0.4$]Refer to caption $\gamma=0.6$]Refer to caption $\gamma=0.8$]Refer to caption
@@ III-A2 Lunar Lander
Figure 5: Success (higher is better), floating, and crash (lower is better) rates for Lunar Lander (top) and Lunar reacher (bottom) with LABEL:sub@fig:$l$-land-assisted-noisy LABEL:sub@fig:$l$-reach-assisted-noisy noisy and LABEL:sub@fig:$l$-land-assisted-laggy LABEL:sub@fig:$l$-reach-assisted-laggy laggy pilots. The dashed blue line denotes the success rate of an expert policy, while the dotted blue line denotes the success rate of our model with full-diffusion ($\gamma=1.0$).
@@ III-A4 Block Pushing
Figure 6: The rates of block pushing task to terminate at correct goal, wrong goal or timeout with LABEL:sub@fig:$l$-reach-assisted-noisy noisy and LABEL:sub@fig:$l$-reach-assisted-laggy laggy pilot. The dashed blue line denotes the rate at which an expert policy reaches the correct goal, while the dotted blue line denotes the same rate for our model with full-diffusion ($\gamma=1.0$).
@@ III-B Training
To understand the pilot’s intent, our copilot relies on a state-conditioned diffusion model $p_{\theta}(a_{t}|s_{t})$. This diffusion model is trained as follows. First, we collect expert demonstrations, each containing a sequence of goal-embedded state-action pairs for each domain, where the goal is randomly sampled at each episode. We then use the resulting demonstrations to train a state-conditioned diffusion model using the DDPM loss (Eqn. 8). The details of the model architecture are described in Appendix A. As we hide a goal from our copilot, we remove goal locations from each observation prior to training.
@@
To collect demonstrations, we first train an expert policy with soft actor-critic [23] for $3\text{M}$ timesteps in Lunar Lander and Lunar Reacher, and $1\text{M}$ timesteps in Block Pushing. We then roll out the policies in each environment to collect demonstrations of state-action pairs $D_{\textrm{expert}}$ for various goals. Appendix A details the model architecture and the hyperparameter settings.
@@ III-D Various pilots with our copilot
Pilot w/o Copilot w/ Copilot w/o Copilot w/ Copilot
Noisy $62.33\pm 2.49$ $\bm{75.67\pm 2.05}$ $\hphantom{0}37.33\pm 2.05$ $\bm{20.00\pm 3.27}$
Laggy $42.00\pm 4.90$ $\bm{74.33\pm 2.87}$ $\hphantom{0}58.00\pm 4.90$ $\bm{21.33\pm 4.50}$
Zero $\hphantom{0}0.00\pm 0.00$ $40.67\pm 2.62$ $100.00\pm 0.00$ $\hphantom{0}6.67\pm 1.70$
Random $\hphantom{0}0.00\pm 0.00$ $16.33\pm 3.09$ $100.00\pm 0.00$ $68.00\pm 1.63$
Expert $99.00\pm 0.82$ $94.67\pm 2.05$ $\hphantom{00}1.00\pm 0.82$ $\hphantom{0}5.33\pm 2.05$
@@ III-F Real Human User Experiments
We conducted a set of human user experiments involving the Lunar Lander and Lunar Reacher continuous control domains. For the experiments, we recruited $17$ participants ($9$ who identified as male, and $8$ as female, with an average age of 2525). [^3] None of the participants were co-authors or otherwise involved in this research. We asked each user to interact with one of two different co-pilots within each episode, one corresponding to direct teleoperation (i.e., no assistance) and the other being our diffusion-based shared autonomy assistant. The identity of the co-pilot was not disclosed to the participant. At the beginning of each experiment, we allowed the user to practice for $10$ episodes with each co-pilot. In the subsequent testing phase, the user controlled the system for another $10$ episodes with each of the co-pilots. We conducted this experiment for both Lunar Lander and Lunar Reacher and emphasize that the location of the goal (the landing pad for Lunar Lander and goal region for Lunar Reacher) was chosen randomly from the continuous space of goals for each episode.
@@
Success Rate $\uparrow$ Crash Rate $\downarrow$ Float Rate Success Rate $\uparrow$ Crash Rate $\downarrow$ Float Rate
w/o Copilot $\hphantom{0}1.76$ $98.24$ $0.00$ $17.06$ $82.94$ $0.00$
w/ Copilot $32.35$ $61.18$ $6.47$ $34.71$ $65.29$ $0.00$
@@ III-G Real-Robot Experiments
We recruited $14$ participants ($10$ who identified as male, $4$ who identified as female, average age of 27.727.7) [^4] None of the participants were co-authors or otherwise involved in this research. for human user experiments with the real robot.
@@
success rate $\uparrow$ episode length $\downarrow$
w/o Copilot $0.89$ $209.23\pm 83.71$
w/ Copilot $1.00$ $143.66\pm 68.42$
@@ Appendix A Model architecture and hyperparameters
We generate the set of demonstrations used to train our diffusion model by collecting 100100k transitions for each goal in the 2D Control environment using the scripted expert policy provided in the original implementation [18], 33M transitions for each goal in the Block Pushing environment using the pretrained expert, and 1010M transitions for Lunar Lander and Lunar Reacher using the corresponding pretrained expert. When we collect expert demonstrations, we filter out episodes that do not succeed, effectively making the quality of the demonstrations better than the original expert. This explains why the assisted expert often performs better than the expert (e.g., Table I—Lunar Reacher).
@@ B-A Lunar Lander and Lunar Reacher
$p_{\text{noise}}=0.2$]Refer to caption $p_{\text{noise}}=0.3$]Refer to caption $p_{\text{noise}}=0.4$]Refer to caption $p_{\text{noise}}=0.5$]Refer to caption
$p_{\text{noise}}=0.6$]Refer to caption $p_{\text{noise}}=0.7$]Refer to caption $p_{\text{noise}}=0.8$]Refer to caption $p_{\text{noise}}=0.9$]Refer to caption
@@
$p_{\text{laggy}}=0.2$]Refer to caption $p_{\text{laggy}}=0.3$]Refer to caption $p_{\text{laggy}}=0.4$]Refer to caption $p_{\text{laggy}}=0.5$]Refer to caption  $p_{\text{laggy}}=0.6$]Refer to caption $p_{\text{laggy}}=0.7$]Refer to caption $p_{\text{laggy}}=0.8$]Refer to caption $p_{\text{laggy}}=0.9$]Refer to caption
@@
$p_{\text{noisy}}=0.2$]Refer to caption $p_{\text{noisy}}=0.3$]Refer to caption $p_{\text{noisy}}=0.4$]Refer to caption $p_{\text{noisy}}=0.5$]Refer to caption  $p_{\text{noisy}}=0.6$]Refer to caption $p_{\text{noisy}}=0.7$]Refer to caption $p_{\text{noisy}}=0.8$]Refer to caption $p_{\text{noisy}}=0.9$]Refer to caption
@@
$p_{\text{laggy}}=0.2$]Refer to caption $p_{\text{laggy}}=0.3$]Refer to caption $p_{\text{laggy}}=0.4$]Refer to caption $p_{\text{laggy}}=0.5$]Refer to caption  $p_{\text{laggy}}=0.6$]Refer to caption $p_{\text{laggy}}=0.7$]Refer to caption $p_{\text{laggy}}=0.8$]Refer to caption $p_{\text{laggy}}=0.9$]Refer to caption
@@
Pilot w/o Copilot ($\gamma=0.0$) w/ Copilot (ours) ($\gamma=0.4$) w/ Copilot ($\gamma=1.0$) w/o Copilot ($\gamma=0.0$) w/ Copilot (ours) ($\gamma=0.4$) w/ Copilot ($\gamma=1.0$)
@@
Noisy $20.67\pm 4.50$ $\bm{68.00\pm 5.35}$ $29.00\pm 2.94$ $\hphantom{0}28.33\pm 2.62$ $\bm{7.67\pm 2.87}$ $18.33\pm 4.92$
Laggy $21.33\pm 2.05$ $\bm{75.00\pm 3.56}$ $26.33\pm 4.78$ $\hphantom{0}76.67\pm 2.49$ $\bm{9.67\pm 3.86}$ $15.67\pm 0.94$
Zero $\hphantom{0}0.00\pm 0.00$ $27.00\pm 0.82$ $28.00\pm 2.16$ $100.00\pm 0.00$ $19.00\pm 2.94$ $13.33\pm 2.49$
Random $\hphantom{0}0.00\pm 0.00$ $25.00\pm 4.32$ $27.67\pm 1.70$ $100.00\pm 0.00$ $19.33\pm 5.25$ $15.67\pm 1.25$
Expert $77.67\pm 2.62$ $78.67\pm 2.87$ $29.33\pm 6.18$ $\hphantom{0}12.33\pm 0.94$ $\hphantom{0}8.00\pm 1.63$ $15.33\pm 1.70$
@@
Noisy $14.33\pm 2.49$ $\bm{45.33\pm 3.30}$ $30.67\pm 4.11$ $\hphantom{0}77.33\pm 3.09$ $\bm{38.00\pm 2.94}$ $47.00\pm 3.56$
Laggy $30.67\pm 5.56$ $\bm{55.33\pm 6.13}$ $30.67\pm 2.62$ $\hphantom{0}69.33\pm 5.56$ $\bm{31.33\pm 2.87}$ $45.67\pm 0.94$
Zero $\hphantom{0}0.00\pm 0.00$ $19.67\pm 2.62$ $31.00\pm 6.98$ $100.00\pm 0.00$ $58.33\pm 3.09$ $45.33\pm 5.56$
Random $\hphantom{0}4.33\pm 1.89$ $22.33\pm 2.87$ $30.00\pm 4.08$ $\hphantom{0}95.33\pm 1.70$ $59.00\pm 0.82$ $45.00\pm 2.16$
Expert $49.33\pm 4.78$ $55.00\pm 2.16$ $28.67\pm 4.50$ $\hphantom{0}44.00\pm 3.56$ $31.67\pm 2.87$ $50.00\pm 1.41$